1.∫(Sin 4 x/Cos 2 x).dx=? 2.∫{1+sin(x/2)} 1/2 dx=? Please give elaborate answers for these at the earliest...

1.∫(Sin4x/Cos2x).dx=?


2.∫{1+sin(x/2)}1/2 dx=?


Please give elaborate answers for these at the earliest...

Grade:12

2 Answers

vikas askiitian expert
13 years ago

I = (sinx/2 + 1)1/2 dx

(sinx/2+1) = (sinx/4)2 + (cosx/4)2 + 2cosx/4sinx/4   =  (sinx/4 + cosx/4)2

after putting this

I ={ sinx/4 + cosx/4 }dx

I= -4cosx/4 + 4sinx/4 + c

please approve my ans if u like it

Aman Bansal
13 years ago

dear chilukuri,

1.∫(Sin4x/Cos2x).dx

Possible intermediate steps:
integral sec(2 x) sin(4 x) dx
Simplify the integrand sin(4 x) sec(2 x) to get 4 sin(x) cos(x):
= integral 4 sin(x) cos(x) dx
Factor out constants:
= 4 integral sin(x) cos(x) dx
For the integrand sin(x) cos(x), substitute u = cos(x) and du = -sin(x) dx:
= -4 integral u du
The integral of u is u^2/2:
= -2 u^2+constant
Substitute back for u = cos(x):
= -2 cos^2(x)+constant

2.∫{1+sin(x/2)}1/2 dx

(4*(-Cos[x/4] + Sin[x/4])*Sqrt[1 + Sin[x/2]])/ (Cos[x/4] + Sin[x/4])

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

AMAN BANSAL